# Topological shape transform for thymus structures

### v1

@@ Introduction
Morphological information has proven invaluable for understanding how structure relates to function in biological tissues, organs, and organisms [1, 2, 3]. Identifying and quantifying distinct shapes are central tasks in biology since geometric differences can indicate variations in developmental processes, disease states, or differences in biological function. The thymus is an essential organ of the adaptive immune system responsible for the maturation, selection, and differentiation of T cells [4]. Within the confines of the thymic microenvironment, T cells are instructed by thymic epithelial cells (TECs) and other stroma cells to establish a diverse, yet self-tolerant T cell receptor (TCR) repertoire - a process known as thymopoiesis [5]. With age, the thymus undergoes involution characterized by reduced total cell numbers, disruption of the epithelial network architecture, and diminished naive T cell output [6, 7]. Paralleling thymic involution and reduced thymopoietic activity, the TCR repertoire becomes constricted with age [8]. Together, this leads to an increased susceptibility to viral infections, heightened autoimmunity, and diminished surveillance of malignant cells due to the immune system’s diminishing capacity to mount effective responses to de-novo antigens [9, 10, 11]. Simultaneously, the thymus undergoes significant architectural remodeling, with both cortical TECs (cTECs) and medullary TECs (mTECs) exhibiting structural changes [2, 12]. To better understand these changes, precise mathematical descriptors are needed to quantify morphological changes at the cell level. A variety of shape descriptors have been developed for biological structures, ranging from deep learning frameworks that leverage large datasets to automatically learn shape representations [13], to techniques grounded in differential geometry that focus on curvature and other geometric invariants [14].
@@
Mouse thymi (n=7) were processed as described in ‘Materials and Methods’ (see below). The boundaries of the thymic epithelial cell (TEC) network were segmented using MATLAB’s Image Processing Toolbox: multilevel thresholding (multithresh) and skeletonization (bwskel) were applied to the cytokeratin 8 (K8) and cytokeratin 14 (K14) TEC markers to generate 2D binary images for all 7 thymi and both markers (values of 1 where the K8/K14 staining is present, 0 where it is absent). A 200 x 200 pixel lattice was superimposed on the 2D images to generate subimages, for subsequent TDA.
@@ Results
To interpret geometric differences between age groups, we use the vectorization of SampEuler to visualize Euler characteristic curves across all directions. We apply the SHAP algorithm to identify which regions of this image are most important for classifying each age group. We repeat training 50 times, retaining trials with accuracy above 50% as valid. For each age group, we average SHAP values across test samples from all valid trials to obtain a final feature importance map (Figure 3D). These maps reveal the distribution of Euler characteristic curves that the classifier recognizes as typical for each age group: regions of high importance indicate where curves consistently lie, effectively summarizing each group’s typical shape signature.
@@ Acknowledgments
For geometric simplicial complex $K$ and $L$ in ${\mathbb{R}}^{d}$, we have:
$\|\text{ECT}(K,v)-\text{ECT}(L,v)\|_{1}\leq\sum_{k=0}^{d-1}2W_{\infty}(\text{Dgm}_{k}(K,v),\text{Dgm}_{k}(L,v)),$
where $\text{Dgm}_{k}(K,v)$ and $\text{Dgm}_{k}(L,v)$ are the $k^{\text{th}}$ persistence diagrams of $K$ and $L$ induced by height filtration along direction $v$.
@@
Let $K,L\subset{\mathbb{R}}^{d}$ be geometric simplicial complexes and let $\varphi:K\rightarrow L$ and $\psi:L\rightarrow K$ be a homotopy equivalence of $M$ and $N$. Let $H_{K}:K\times I\rightarrow K$ and $H_{L}:L\times I\rightarrow L$ be homotopies from $\psi\circ\varphi$ to $\text{Id}_{K}$ and from $\varphi\circ\psi$ to $\text{Id}_{L}$ respectively. We denote by $\|\cdot\|_{{\mathbb{R}}^{d}}^{2}$ the normal Euclidean 2-norm. If there exists $\epsilon$ strictly positive such that $\|x-\varphi(x)\|^{2}_{{\mathbb{R}}^{d}}\leq\epsilon$ and $\|y-\psi(y)\|^{2}_{{\mathbb{R}}^{d}}\leq\epsilon$ for all $x\in K$ and $y\in L$, as well as $\|x-H_{K}(x,s)\|^{2}_{{\mathbb{R}}_{d}}\leq 2\epsilon$ and $\|y-H_{L}(y,s)\|^{2}_{{\mathbb{R}}_{d}}\leq 2\epsilon$ for all $x\in K,y\in L,s\in I$, then we say $K$ and $L$ are $\epsilon$-controlled homotopic.
@@
$d_{\text{ECT}}^{1}(K,L)\leq 2d\epsilon.$
@@
$\leq\int_{S^{d-1}}\sum_{k=0}^{d-1}2W_{\infty}(\text{Dgm}_{k}(K,v),\text{Dgm}_{k}(L,v))\,\mathrm{d}v$
$=\sum_{k=0}^{d-1}2\int_{S^{d-1}}W_{\infty}(\text{Dgm}_{k}(K,v),\text{Dgm}_{k}(L,v))\,\mathrm{d}v$
$\leq\sum_{k=0}^{d-1}2\max_{v\in S^{d-1}}W_{\infty}(\text{Dgm}_{k}(K,v),\text{Dgm}_{k}(L,v))$
$=2\sum_{k=0}^{d-1}d_{\infty}^{\text{PHT}^{k}}(K,L)\leq 2d\epsilon.$
@@
The ECT pushforward measures equipped with the Wasserstein distance $W_{1}$ is Lipschitz continuous for homotopic equivalent complexes.
@@
This theorem not only shows that the pushforward measure is stable under Wasserstein distance but also more stable than normal ECT outputs, as it always less than or equal to the minimal distance between two shapes up to $O(d)$ actions.
@@
We will make use of the following lemma in the proof, proving in [22] along the proof of Theorem S3.
@@
In this section, we show how vectorization of SampEuler vectorization provides interpretable topological features. The vectorization creates an intensity map of ECC curves across all directions, revealing typical topological patterns for a given shape. We demonstrate how these patterns can be interpreted in terms of the underlying geometric structure.
@@
Figure S3: A: i) 50 geometric simplicial complex representations of ellipses with centres uniformly sampled from the centred square of side length 50. ii) The corresponding vectorized SampEuler result. B: i) 40 geometric simplicial complex representations of ellipses with centres uniformly sampled from the centred square of side length 50. ii) The corresponding vectorized SampEuler result. C: 50 geometric simplicial complex representations of ellipses with centres uniformly sampled from the restriction of quadrant 2, 3, 4 of the centred square of side length 50. ii) The corresponding vectorized SampEuler result. D: 50 geometric simplicial complex representations of ellipses with centres uniformly sampled from a centred ellipse with major axis length 100 and minor axis length 20. ii) The corresponding vectorized SampEuler result.

### v2

@@ Introduction
Morphological information has proven invaluable for understanding how structure relates to function in biological tissues, organs, and organisms [1, 2, 3]. Identifying and quantifying distinct shapes are central tasks in biology since geometric differences can indicate variations in developmental processes, disease states, or differences in biological function. The thymus is an essential organ of the adaptive immune system responsible for the maturation, selection, and differentiation of T cells [4]. Within the confines of the thymic microenvironment, T cells are instructed by thymic epithelial cells (TECs) and other stromal cells to establish a diverse, yet self-tolerant T cell receptor (TCR) repertoire - a process known as thymopoiesis [5]. With age, the thymus undergoes involution characterized by reduced total cell numbers, disruption of the epithelial network architecture, and diminished naive T cell output [6, 7]. Paralleling thymic involution and reduced thymopoietic activity, the TCR repertoire becomes constricted with age [8]. Together, this leads to an increased susceptibility to viral infections, heightened autoimmunity, and diminished surveillance of malignant cells due to the immune system’s diminishing capacity to mount effective responses to de-novo antigens [9, 10, 11]. Simultaneously, the thymus undergoes significant architectural remodeling, with both cortical TECs (cTECs) and medullary TECs (mTECs) exhibiting structural changes [2, 12]. To better understand these changes, precise mathematical descriptors are needed to quantify morphological changes at the cell level. A variety of shape descriptors have been developed for biological structures, ranging from deep learning frameworks that leverage large datasets to automatically learn shape representations [13], to techniques grounded in differential geometry that focus on curvature and other geometric invariants [14].
@@
Mouse thymi (n=7) were processed as described in “Materials and Methods” (see below). The boundaries of the thymic epithelial cell (TEC) network were segmented using MATLAB’s Image Processing Toolbox: multilevel thresholding (multithresh) and skeletonization (bwskel) were applied to the cytokeratin 8 (K8) and cytokeratin 14 (K14) TEC markers to generate 2D binary images for all 7 thymi and both markers (values of 1 where the K8/K14 staining is present, 0 where it is absent). A 200 x 200 pixel lattice was superimposed on the 2D images to generate subimages, for subsequent TDA.
@@ Results
To interpret geometric differences between age groups, we use the vectorization of SampEuler to visualize Euler characteristic curves across all directions. We apply the SHAP algorithm to identify which regions of this image are most important for classifying each age group. We repeat training 50 times, retaining trials with accuracy above 80% as valid. For each age group, we average SHAP values across test samples from all valid trials to obtain a final feature importance map (Figure 3D). These maps reveal the distribution of Euler characteristic curves that the classifier recognizes as typical for each age group: regions of high importance indicate where curves consistently lie, effectively summarizing each group’s typical shape signature.
@@ Acknowledgments
For geometric simplicial complexes $K$ and $L$ in ${\mathbb{R}}^{d}$, we have:
$\|\text{ECT}(K,v)-\text{ECT}(L,v)\|_{1}\leq\sum_{k=0}^{d-1}2n_{k}W_{\infty}(\text{Dgm}_{k}(K,v),\text{Dgm}_{k}(L,v)),$
where $\text{Dgm}_{k}(K,v)$ and $\text{Dgm}_{k}(L,v)$ are the $k^{\text{th}}$ persistence diagrams of $K$ and $L$ induced by height filtration along direction $v$, and $n_{k}$ is the maximum number of points in $\text{Dgm}_{k}(K,v)$ or $\text{Dgm}_{k}(L,v)$.
@@
Let $K,L\subset{\mathbb{R}}^{d}$ be geometric simplicial complexes and let $\varphi:K\rightarrow L$ and $\psi:L\rightarrow K$ be a homotopy equivalence of $K$ and $L$. Let $H_{K}:K\times I\rightarrow K$ and $H_{L}:L\times I\rightarrow L$ be homotopies from $\psi\circ\varphi$ to $\text{Id}_{K}$ and from $\varphi\circ\psi$ to $\text{Id}_{L}$ respectively. We denote by $\|\cdot\|_{{\mathbb{R}}^{d}}^{2}$ the normal Euclidean 2-norm. If there exists $\epsilon$ strictly positive such that $\|x-\varphi(x)\|^{2}_{{\mathbb{R}}^{d}}\leq\epsilon$ and $\|y-\psi(y)\|^{2}_{{\mathbb{R}}^{d}}\leq\epsilon$ for all $x\in K$ and $y\in L$, as well as $\|x-H_{K}(x,s)\|^{2}_{{\mathbb{R}}_{d}}\leq 2\epsilon$ and $\|y-H_{L}(y,s)\|^{2}_{{\mathbb{R}}_{d}}\leq 2\epsilon$ for all $x\in K,y\in L,s\in I$, then we say $K$ and $L$ are $\epsilon$-controlled homotopic.
@@
Let $K,L\subset{\mathbb{R}}^{d}$ be geometric simplicial complexes that are $\epsilon$-controlled homotopic, and let $n_{k}$ be the maximum number of bars in $\text{Dgm}_{k}(K,v)$ or $\text{Dgm}_{k}(L,v)$ for all $v\in S^{d-1}$, and let $n=\sum_{k=0}^{d-1}n_{k}$. Then we have:
$d_{\text{ECT}}^{1}(K,L)\leq 2dn\epsilon.$
@@
$\leq\int_{S^{d-1}}\sum_{k=0}^{d-1}2n_{k}W_{\infty}(\text{Dgm}_{k}(K,v),\text{Dgm}_{k}(L,v))\,\mathrm{d}v$
$=\sum_{k=0}^{d-1}2n_{k}\int_{S^{d-1}}W_{\infty}(\text{Dgm}_{k}(K,v),\text{Dgm}_{k}(L,v))\,\mathrm{d}v$
$\leq\sum_{k=0}^{d-1}2n_{k}\max_{v\in S^{d-1}}W_{\infty}(\text{Dgm}_{k}(K,v),\text{Dgm}_{k}(L,v))$
$=2\sum_{k=0}^{d-1}n_{k}d_{\infty}^{\text{PHT}^{k}}(K,L)\leq 2dn\epsilon.$
@@
The ECT pushforward measure equipped with the Wasserstein distance $W_{1}$ is Lipschitz continuous for homotopic equivalent complexes.
@@
This theorem not only shows that the pushforward measure is stable under Wasserstein distance but also more stable than normal ECT outputs, as it is always less than or equal to the minimal distance between two shapes up to $O(d)$ actions.
@@
We will make use of the following lemma in the proof, proven in [22] along the proof of Theorem S3.
@@
In this section, we show how vectorization of SampEuler provides interpretable topological features. The vectorization creates an intensity map of ECC curves across all directions, revealing typical topological patterns for a given shape. We demonstrate how these patterns can be interpreted in terms of the underlying geometric structure.
@@
Figure S3: A: i) 50 geometric simplicial complex representations of ellipses with centres uniformly sampled from the centred square of side length 50. ii) The corresponding vectorized SampEuler result. B: i) 40 geometric simplicial complex representations of ellipses with centres uniformly sampled from the centred square of side length 50. ii) The corresponding vectorized SampEuler result. C: i) 50 geometric simplicial complex representations of ellipses with centres uniformly sampled from the restriction of quadrant 2, 3, 4 of the centred square of side length 50. ii) The corresponding vectorized SampEuler result. D: 50 geometric simplicial complex representations of ellipses with centres uniformly sampled from a centred ellipse with major axis length 100 and minor axis length 20. ii) The corresponding vectorized SampEuler result.